# The Automated but Risky Game: Modeling Agent-to-Agent Negotiations and Transactions in Consumer Markets

### v3

@@ Appendix C Discussion
Appendix D Details of Dataset
D.1 Data Structure
@@
"Reference": "https://www.toyota.com/camry/"
@@
D.2 Wholesale Generation Prompt
@@
Appendix E Details of Metrics.
E.1 Main
Price Reduction Rate(PRR). The Price Reduction Rate(PRR) quantifies the relative price change achieved through negotiation:
@@ E.1 Main
A higher BR indicates stronger buyer bargaining power, while the seller concedes more, reflecting weaker negotiation strength.
Relative Profit (RP). We define the Relative Profit (RP) as the ratio between the total profit achieved by the model and the minimum reference profit (e.g. the GPT-3.5 profit in main experiment):
@@
Deal Rate(DR). The Deal Rate (DR) measures the percentage of negotiations that result in a successful transaction:
@@
[…]
Profit Rate (PR). We define the Profit Rate (PR) as the average per-product profit margin across all successful transactions. For each deal, the profit margin is computed relative to the wholesale cost. Formally:
@@
E.2 Anomaly
Out of Budget Rate (OBR). The Out of Budget Rate (OBR) quantifies how often the final accepted price exceeds the buyer’s budget constraint:
@@ E.2 Anomaly
Out of Wholesale Rate (OWR). The Out of Wholesale Rate (OWR) measures how often the final accepted price falls below the wholesale price, indicating unprofitable transactions from the seller’s perspective:
@@
Overpayment Rate (OPR). The Overpayment Rate (OPR) quantifies how often the buyer ends up paying more than the reference retail price of the product in a successful transaction:
@@
Deadlock Rate (DLR). The Deadlock Rate (DLR) quantifies the proportion of negotiations that reach the maximum allowed number of rounds $T_{\max}$ without reaching any agreement:
@@
Metric Definition and Description
Total Profit Cumulative profit across all successful negotiations:
$TP=\sum_{i=1}^{N_{\text{deal}}}(p_{a}^{T,(i)}-p_{w}^{(i)})$
Relative Profit Ratio of current model’s profit to the worst-performing model’s profit:
$RP=\frac{TP}{TP_{\text{min}}}$
Profit Rate Average profit margin relative to wholesale price over successful deals:
$PR=\frac{1}{N_{\text{deal}}}\sum_{i=1}^{N_{\text{deal}}}\frac{p_{a}^{T,(i)}-p_{w}^{(i)}}{p_{w}^{(i)}}$
Out of Budget Rate Fraction of negotiations where final price exceeds buyer’s fixed budget $\beta$:
$OBR=\frac{N_{\text{over}}}{N}$, where $p_{a}^{T,(i)}>\beta$
Out of Wholesale Rate Fraction of negotiations where final price falls below the wholesale price:
$OWR=\frac{N_{\text{below}}}{N}$, where $p_{a}^{T,(i)}<p_{w}^{(i)}$
Overpayment Rate Fraction of successful deals where buyer pays more than the retail price:
$OPR=\frac{N_{\text{over}}}{N}$, where $p_{a}^{T,(i)}>p_{r}^{(i)}$
Deadlock Rate Fraction of negotiations that reach the maximum round limit $T_{\max}$ without any agreement:
$DR=\frac{N_{\text{deadlock}}}{N}$
Table 4: Summary of Evaluation Metrics
Appendix F Details of Negotiation Implementation
F.1 System Prompt of Buyer
@@ F.2 Greeting Prompt
F.3 System Prompt of Seller
@@
F.4 System Prompt of Judge
@@
F.5 System Prompt of Analyst
@@
Appendix G Details of More Results
G.1 Understanding the Negotiation Gap via Model Specifications and Common Benchmarks.
To investigate the sources of variation in negotiation capacity across models, we collect data on four commonly referenced model characteristics as potential explanatory factors. [^2] We obtain these data from model providers’ official websites or technical papers: https://openai.com/index/hello-gpt-4o/; https://arxiv.org/abs/2501.12948; https://qwenlm.github.io/blog/qwen2.5-llm/. The parameter count for GPT-4o-mini is estimated based on analysis in https://arxiv.org/abs/2412.19260. , including one architectural attribute: model size (in billions of parameters), and three performance-based benchmarks: general task performance (MMLU (hendrycks2020measuring)), mathematical ability (MATH (hendrycksmath2021)), and scientific ability (GPQA (rein2024gpqa)). We combine three negotiation-relevant metrics—Buyer Price Reduction Rate($PRR_{\text{Buyer}}$), reverse of Seller Price Reduction($1-PRR_{\text{Seller}}$), and $RP$—into a scalar indicator via z-score normalization followed by averaging, yielding a composite Negotiation Capacity Score ($NCS$). We then compute the Pearson correlation between each model’s $NCS$ and the four benchmark scores. As shown in Figure 18, negotiation capacity shows a very strong correlation with general task performance on MMLU ($r=0.93$), along with substantial correlations with mathematical ($r=0.87$) and scientific ability ($r=0.80$). The weakest correlation appears with model size ($r=0.53$), which we attribute to multiple factors: some high-parameter models (e.g., GPT-3.5) belong to earlier generations with less optimized architectures and performance, while for commercial models such as GPT-4o-mini, exact parameter counts are unavailable and must be estimated from external sources.
G.2 Negotiation Capacity Gap Indicates Behavioral Robustness Gap.
@@ Appendix G Details of More Results
Figures 6, 7, and 8 present anomaly indicators across six models analyzed in Section 3.2. The data reveals a notable pattern: the proportion of anomalies appears inversely related to the models’ negotiation capabilities. This observation motivates the research question: Are models with stronger negotiation skills also more robust against automation-induced anomalies?
@@
To investigate this relationship, we reuse the previously defined Negotiation Capacity Score (NCS) (see Section G.1). To quantify a model’s overall tendency toward negotiation anomalies, we construct a composite Risk Index by aggregating the four anomaly-related indicators introduced in Section 4.2. Each indicator is standardized using z-score normalization and averaged to produce a unified scalar value. We then compute the Pearson correlation between NCS and the Risk Index. As shown in Figure 19, the result ($r=-0.67$) indicates a moderate negative association: models with higher negotiation capacity consistently exhibit lower anomaly indices, suggesting greater behavioral robustness in automated negotiation systems.

### v4

@@ Appendix B Discussion
Appendix C Details of Dataset
C.1 Data Structure
@@
"Reference": "https://www.toyota.com
/camry/"
@@
C.2 Wholesale Generation Prompt
@@
Appendix D Details of Metrics.
D.1 Main
Price Reduction Rate(PRR).
The Price Reduction Rate(PRR) quantifies the relative price change achieved through negotiation:
@@ D.1 Main
A higher PRR indicates stronger buyer bargaining power, while the seller concedes more, reflecting weaker negotiation strength.
Relative Profit (RP).
We define the Relative Profit (RP) as the ratio between the total profit achieved by the model and the minimum reference profit (e.g. the GPT-3.5 profit in main experiment):
@@
Deal Rate(DR).
The Deal Rate (DR) measures the percentage of negotiations that result in a successful transaction:
@@
In here, $|N_{\text{deal}}|$ is the number of successful negotiations. $|N$ is the total number of negotiations.
Profit Rate (PR).
We define the Profit Rate (PR) as the average per-product profit margin across all successful transactions. For each deal, the profit margin is computed relative to the wholesale cost. Formally:
@@
D.2 Anomaly
Out of Budget Rate (OBR).
The Out of Budget Rate (OBR) quantifies how often the final accepted price exceeds the buyer’s budget constraint:
@@ D.2 Anomaly
Out of Wholesale Rate (OWR).
The Out of Wholesale Rate (OWR) measures how often the final accepted price falls below the wholesale price, indicating unprofitable transactions from the seller’s perspective:
@@
Overpayment Rate (OPR).
The Overpayment Rate (OPR) quantifies how often the buyer ends up paying more than the reference retail price of the product in a successful transaction:
@@
Deadlock Rate (DLR).
The Deadlock Rate (DLR) quantifies the proportion of negotiations that reach the maximum allowed number of rounds $T_{\max}$ without reaching any agreement:
@@
Appendix E Details of Negotiation Implementation
E.1 System Prompt of Buyer
@@ E.2 Greeting Prompt
E.3 System Prompt of Seller
@@
E.4 System Prompt of Judge
@@
E.5 System Prompt of Analyst
@@
Appendix F Details of More Results
F.1 Negotiation Capacity Gap Indicates Behavioral Robustness Gap.
@@ Appendix F Details of More Results
Figures 6, 8, and 7 present anomaly indicators across six models analyzed in Section 3.2. The data reveals a notable pattern: the proportion of anomalies appears inversely related to the models’ negotiation capabilities. This observation motivates the research question: Are models with stronger negotiation skills also more robust against automation-induced anomalies?
@@
To investigate this relationship, we reuse the previously defined Negotiation Capacity Score (NCS) (see Section 3.4). To quantify a model’s overall tendency toward negotiation anomalies, we construct a composite Risk Index by aggregating the four anomaly-related indicators introduced in Section 4.1. Each indicator is standardized using z-score normalization and averaged to produce a unified scalar value. We then compute the Pearson correlation between NCS and the Risk Index. As shown in Figure 19, the result (r=0.67r=-0.67) indicates a moderate negative association: models with higher negotiation capacity consistently exhibit lower anomaly indices, suggesting greater behavioral robustness in automated negotiation systems.
F.2 From Model Capability Gap to Economic Loss
In Sections §3.2, we discuss the capability gap of different models and also the asymmetric influence of buyer versus seller agent roles. Although such performance gaps may seem expected in experiments, deploying such agents in consumer settings could systematically disadvantage users who rely on less capable models.
In particular, we view these interactions as imbalanced games, where one party deploys a significantly stronger agent than the other. Whether a strong buyer faces a weak seller or vice versa, the party with the weaker agent suffers a strategic disadvantage. Thus, one crucial question emerges: How does this strategic disadvantage translate into quantifiable economic loss?
Buyer Seller Avg Payment(\$) $\boldsymbol{\Delta}$ from Baseline (%) Impact
Strong vs. Strong
DeepSeek-R1 DeepSeek-R1 1,423,090 Baseline
Weak-Buyer vs. Strong-Seller
GPT-3.5 DeepSeek-R1 1,452,699 +2.09% Buyer overpays by 2.09%
Qwen-7B DeepSeek-R1 1,454,633 +2.09% Buyer overpays by 2.09%
Qwen-14B DeepSeek-R1 1,438,834 +1.10% Buyer overpays by 1.10%
Strong-Buyer vs. Weak-Seller
DeepSeek-R1 GPT-3.5 1,221,980 –14.13% Seller earns 14.13% less
DeepSeek-R1 Qwen-7B 1,314,796 –7.62% Seller earns 7.62% less
DeepSeek-R1 Qwen-14B 1,325,570 –6.94% Seller earns 6.94% less
Asymmetric Influence of Agent Roles.
As shown in Figure 20 (top), the heatmap illustrates the $PRR$ across all pairwise combinations of buyer and seller agents. Our analysis reveals a clear asymmetry in agent roles: the choice of the seller model has a significantly larger impact on negotiation outcomes than the choice of the buyer model. For example, when we fix the seller as GPT-3.5 and vary the buyer agents, the difference between the highest and lowest $PRR$ is only 2.6%. In contrast, when we fix the buyer as GPT-3.5 and vary the seller agents, the $PRR$ gap reaches up to 14.9%. This asymmetry also explains the observation in Figure 2 (top), where the average $PRR$ across different buyer agents shows relatively small variance: buyers have less influence on the final negotiation result compared to sellers.
Figure 20: Top: Average PRRPRR heatmaps over 5 budget settings per agent pair; Bottom: Average Deal Rate of seller agents over 5 budgets settings.
Budget as a Window into Seller Strategy.
Does the buyer’s budget affect the seller’s strategy? From Figure 20 (bottom), models such as GPT-4.1, o4-mini, and DeepSeek-R1—identified as the most profitable sellers, demonstrates adaptability across various budget scenarios without explicit budget knowledge by effectively adjusting deal rates based on negotiation dynamics. Conversely, GPT-4o-mini and o3 consistently underperforms with below-average deal rates across all budget levels. Low transaction volume undermines total revenue despite any profit margin advantages. GPT-3.5 and Qwen2.5-7b maintains above-average deal rates in all settings, potentially indicating aggressive pricing strategies that secure deals but yield lower profit margins.
Appendix G Details of RL Training
We formulate prompt optimization as an online multi-armed bandit problem with $K=96$ candidate prompts, each corresponding to a distinct negotiation strategy configuration. The training process proceeds as follows.
Core Policy Update.
At step $t$, the policy over $K$ actions is defined by a softmax distribution:
$\pi_{i}(\theta)$ $=\frac{e^{\theta_{i}}}{\sum_{j=1}^{K}e^{\theta_{j}}},\quad i=1,\dots,K,$
$a_{t}$ $\sim\pi(\theta),\quad r_{t}\in\mathbb{R}.$
We maintain an exponential moving average baseline
$b_{t}$ $=0.9\,b_{t-1}+0.1\,r_{t},$
$A_{t}$ $=r_{t}-b_{t},$
where $A_{t}$ is the advantage. The update rule modifies only the chosen action:
$\theta_{a_{t}}\leftarrow\theta_{a_{t}}+\eta\,A_{t}\,\bigl(1-\pi_{a_{t}}\bigr).$
Reward Shaping.
The reward $r_{t}$ penalizes negotiation anomalies according to budget conditions:
  • High budget with overpayment: 2.0-2.0
  • Low budget with out-of-budget violation: 1.0-1.0
  • Negotiation deadlock: 1.0-1.0
Prompt Action Space.
We instantiate the bandit arms as a combinatorial prompt space $\mathcal{A}$ of size $\lvert\mathcal{A}\rvert=96$, formed by the Cartesian product of several independent design axes with all other fields fixed for the first batch (e.g., refusal_tone=polite, brevity=short, self_check_clause=strict). The main axes are:
  • Budget emphasis (2): {hard, medium_hard}.
  • Price increase policy (2): {end_now, warn_then_end}.
  • Exit turns under no progress (3): {2, 3, 4}.
  • Progress threshold (2): {tiny=0.3%, small=0.8%}.
  • Concession style (2): {none, tiny_steps}.
  • Non-price ask (2): {False, True}.
Training Workflow.
We adopt the following schedule:
  • Warmup coverage. Each action is sampled once under both high- and low-budget conditions.
  • Main training. The main training phase proceeds as follows. First, budget sampling is scheduled such that the first half of training uses only low-budget settings, while the second half samples high budgets with probability $0.7$. Second, exploration annealing is applied, where the $\varepsilon$-greedy rate decays linearly from $0.10$ to $0.02$ over global training progress. Third, an active set restriction is enforced: training begins with $K=24$ active actions and shrinks to $K=12$ after two-thirds of progress, with sampling restricted to this set using the normalized distribution $\pi(\theta)$. Finally, the sampling policy mixes $\varepsilon$-random exploration with softmax sampling: at each step, with probability $\varepsilon$ a random action from the active set is chosen; otherwise, actions are drawn according to $\pi(\theta)$. Every $N$ steps, the least-sampled action in the active set is forcibly selected to ensure coverage.
Output.
At the end of training, the best single prompt is selected as
$a^{\star}=\arg\max_{i\in\{1,\dots,K\}}\theta_{i}.$